# Improving Entropy-Based Test-Time Adaptation from a Clustering View

### v5

@@ Conclusion
In this paper, we revisited existing EBTTA methods from the view of clustering. We found that the forward process can be interpreted as the label assignment while the backward process of EBTTA can be interpreted as the update of centers in clustering. This view can guide us to improve EBTTA by overcoming weaknesses in clustering methods, which are sensitive to initial assignments, nearest neighbor information, outliers, and batch size. We proposed to assign robust labels, use spectral embedding, select samples and accumulate gradients. Conducted experimental results have demonstrated the superiority of our proposed method. In our future work, we wish to give a better theoretical explanation for more test-time adaptation methods.

### v6

@@ Conclusion
In this paper, we revisited existing EBTTA methods from the view of clustering. We found that the forward process can be interpreted as the label assignment while the backward process of EBTTA can be interpreted as the update of centers in clustering. This view can guide us to improve EBTTA by overcoming weaknesses in clustering methods. We proposed to assign robust labels, use spectral embedding, select samples and accumulate gradients. Conducted experimental results have demonstrated the superiority of our proposed method. In our future work, we wish to give a better theoretical explanation for more test-time adaptation methods.